If 2x+1/3x=6 then 3x+1/2x is equal to?
Answer by checkley79(3341)

2x+1/3x=6
(2 1/3)X=6
(7/3)X=6
7X/3=6 CROSS MULTIPLY
7X=6*3
7X=18
X=18/7 ANS.
################
3x+1/2x
3*18/7+(1/2)(18/7)
54/7+18/14
54/7+9/7=63/7=9 ANS.